Vargas Corporation is working on its direct labor budget for the next two months. Each unit of output requires 0.88 direct labor-hours. The direct labor rate is $12.30 per direct labor-hour. The production budget calls for producing 8,200 units in October and 8,000 units in November. The company guarantees its direct labor workers a 40-hour paid work week. With the number of workers currently employed, that means that the company is committed to paying its direct labor work force for at least 7,250 hours in total each month even if there is not enough work to keep them busy. What would be the total combined direct labor cost for the two months?

Multiple Choice

  • $178,350.00.

  • $175,348.80.

  • $175,767.00.

  • $177,931.80.
Answer #1

Labor hours needed in October = 8,200*0.88 = 7216

Labor hours needed in November = 8,000*0.88 = 7040

As the hours needed in both months are below 7,250 hours (min)

Combined direct label cost

= (7250*2 months) *12.30

= 178,350